# Universal finite-size scaling in high-dimensional critical phenomena

### v2

@@ I Introduction
[…]
[scale = 0.6]tree_torus
[…]
[scale = 0.6]tree_lift2
@@
Beyond the computation of FSS critical exponents, more precise information about the scaling of the susceptibility and two-point function plateau can be sought. For the $n$-component $|\varphi|^{4}$ model in dimensions $d>4$ with PBC, an exact profile for the amplitude of the susceptibility in the window was computed in 1985 using a Wilsonian renormalisation group method [28, 29]; we review this computation in Appendix D. A mathematically rigorous derivation of the same profile has been carried out recently on the hierarchical lattice [30, 31], for dimensions $d\geq 4$ (including $d=d_{\mathrm{c}}=4$). For $n=1$, the profile is the same as its analogue on the complete graph (Curie–Weiss model). Despite the results of [28], the profile has subsequently received scant attention in the FSS literature, for example it is not explicitly mentioned in any of [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. Part of our purpose is to focus attention on the importance of the profile. The profiles for self-avoiding walk and branched polymers on the complete graph have been computed, and we conjecture that these profiles also apply to SR and LR models at and above their critical dimensions. For percolation, we also present a conjecture for the profile, which is of a different character than the others. Our conjectures for the universal profiles are stated in Section VI.
@@ III Proof of plateau theorem
$c_{0}\frac{\chi(\beta_{*})^{d_{\mathrm{c}}/2}}{V}\leq c_{0}^{\prime}\frac{1}{Vt_{*}^{\gamma d_{\mathrm{c}}/2}}=\frac{1}{c_{1}^{\gamma d_{\mathrm{c}}/2}}\leq\frac{1}{2}.$ (17)
@@ IV Long-range self-avoiding walk
[…]
@@ V Further applications of the plateau theorem
The plateau theorem has been isolated here, and also applied to long-range models, for the first time. However, it has previously been used implicitly in multiple contexts. We now summarise the relevant literature; the conclusions are presented in Table 1. (Logarithmic corrections at the critical dimension are discussed in [30, 31, 35, 36] for SR models.)
@@
The results concern the following kernels $J$ on $\mathbb{Z}^{d}$ and $\bar{J}_{R}$ on $\mathbb{T}_{R}^{d}$ for short-range and long-range models:
@@
  • spread-out LR: $J_{0,x}\asymp L^{-d}(1+|x|/L)^{-(d+\alpha)}$ with $\alpha\in(0,2)$ for $d>1$ and $\alpha\in(0,1)$ for $d=1$ ($L$ is fixed and large).
@@
  • For the SR $1$-component $\varphi^{4}$ model, (6)–(7) are proved in [39]. For the torus, although Theorem 1 has not yet been proved, much stronger results have been proved on the hierarchical lattice for $d\geq 4$ [30, 31], as discussed in more detail in Section VI.
@@ VI Universal profiles
which extends the definition of $f_{n}$ to all n>2n>-2. The $n=0$ case reduces to $f_{0}(s)=I_{1}(s)$ (since $I_{3}(x)+sI_{1}(s)=1$) as in Table 2; this is the profile for SAW on the complete graph [30, Appendix B] and is consistent with SAW being the $n=0$ spin model [43]. The limiting case $f_{-2}(s)=s^{-1}$ is the correct profile for a Gaussian model. Percolation on the complete graph is the Erdős–Rényi random graph. Based on known results on the distribution of the Erdős–Rényi clusters, its profile is conjectured in [17, Appendix C] to be the above function $f_{\mathrm{perc}}$.
@@ VII Free boundary conditions
$\chi_{R}^{\mathrm{F}}({\beta_{\mathrm{c}}})\asymp R^{\alpha},\qquad G_{R,{\beta_{\mathrm{c}}}}^{\mathrm{F}}(0,x)\asymp\frac{1}{|x|^{d-\alpha}}.$ (47)
(Since translation invariance is broken, we write both points in $G_{R,{\beta_{\mathrm{c}}}}^{\mathrm{F}}(0,x)$; the origin is the centre of the box.) In contrast with (4) for PBC, there is no plateau in (47). This has been proved for the SR Ising model [44] (with the hypothesis supplied by [39, 22]) and SR percolation [45].
@@
Since $d>d_{\mathrm{c},\alpha}=\frac{\alpha}{2}d_{\mathrm{c}}$, the shift $R^{-\frac{1}{\nu}}=V^{-\frac{\alpha}{\gamma d}}$ is larger than the window width $V^{-\frac{2}{\gamma d_{\mathrm{c}}}}$, and the PBC and FBC windows do not overlap. This shifting of the PBC scaling behaviour to the FBC pseudocritical point was observed numerically for the 5D SR Ising model and is discussed in [9, 4] (contrary to the conclusions of [13]). In Appendix D, we provide strong evidence in favour of Conjecture 6.
@@ Acknowledgements
We thank our collaborators on aspects of this work: Tom Hutchcroft, Emmanuel Michta, Romain Panis. The work of YL and GS was supported in part by NSERC of Canada. YL and GS thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme Stochastic systems for anomalous diffusion, where work on this paper was undertaken; this work was supported by EPSRC grant EP/R014604/1. The work of GS was supported in part by the Clay Mathematics Institute. JP thanks the Seoul National University for hospitality during a visit when this work was partially undertaken.
@@ Appendix A Elementary lemmas
If the maximum in the lower summation limit is 11 then $\xi/R\leq 1$ and ($5$) is of order $(\xi/R)^{a+\varepsilon}\leq(\xi/R)^{a}$. If instead the maximum is $\lceil 2\xi/R\rceil$ then (55) is bounded by a multiple of $(\xi/R)^{a+\varepsilon}(\xi/R)^{-\varepsilon}=(\xi/R)^{a}$. This completes the proof. ∎
@@ B.4 Ising model
For the SR $n$-component $|\varphi|^{4}$ model in the critical dimension $d=d_{\mathrm{c}}=4$, it is well-known and rigorously proved [50, 51] that $\gamma=2\nu=1$ with logarithmic exponents $\hat{\gamma}=2\hat{\nu}=\frac{n+2}{n+8}$. Also, the specific heat has exponent $\alpha=0$ for $n\geq 1$, with $\hat{\alpha}=\frac{4-n}{n+8}$ for $n=1,2,3$, whereas for $n=4$ the specific heat diverges as $\log\log t^{-1}$ and for $n\geq 4$ it does not diverge at all [50]. In particular, $\hat{\alpha}$ is never strictly negative. Let $\hat{\theta}=\frac{4-n}{2(n+8)}=\frac{1}{2}-\hat{\gamma}$, for all $n\geq 1$ including $n\geq 4$. It is argued in [35, (3.6), (4.3)], and proved rigorously for the hierarchical model in [30], that under PBC the finite-volume susceptibility has a log correction $\chi_{R}\asymp V^{1/2}(\log V)^{1/2}$ and the window width is $V^{-1/2}(\log V)^{-\hat{\theta}}$. At the high-temperature edge of the window, the infinite-volume correlation length therefore obeys
@@ Appendix D Renormalisation group
Universal ratios can be computed similarly. E.g., with the same $\nu=\nu_{\mathrm{c}}+sa_{d}^{-1}\bar{g}^{1/2}V^{-1/2}$, the moments of the total spin $\Phi_{R}=\sum_{x}\varphi_{x\in\mathbb{T}_{R}^{d}}$ obey, as $R\to\infty$,

### v3

@@ I Introduction
Beyond the computation of FSS critical exponents, more precise information about the scaling of the susceptibility and two-point function plateau can be sought. For the $n$-component $|\varphi|^{4}$ model in dimensions $d>4$ with PBC, an exact profile for the amplitude of the susceptibility in the window was computed in 1985 using a Wilsonian renormalisation group method [28, 29]; we review this computation in Appendix D. A mathematically rigorous derivation of the same profile has been carried out recently on the hierarchical lattice [30, 31] for dimensions $d\geq 4$ (including $d=d_{\mathrm{c}}=4$). For $n=1$, the profile is the same as its analogue on the complete graph (Curie–Weiss model). Despite the results of [28], the profile has subsequently received scant attention in the FSS literature, for example it is not explicitly mentioned in any of [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. Part of our purpose is to focus attention on the importance of the profile. The profiles for self-avoiding walk and branched polymers on the complete graph have been computed, and we conjecture that these profiles also apply to SR and LR models at and above their critical dimensions. We also discuss a profile for percolation, which is of a different character than the others. Our conjectures for the universal profiles are stated in Section VI.
@@ III Proof of plateau theorem
$c_{0}\frac{\chi(\beta_{*})^{d_{\mathrm{c}}/2}}{V}\leq c_{0}^{\prime}\frac{1}{Vt_{*}^{\gamma d_{\mathrm{c}}/2}}=\frac{c_{0}^{\prime}}{c_{1}^{\gamma d_{\mathrm{c}}/2}}\leq\frac{1}{2}.$ (17)
@@ IV Long-range self-avoiding walk
To begin, we observe from (IV) that the difference $\Gamma_{R,\beta}-G_{R,\beta}$ arises from self-avoiding walks $\omega$ whose projection $\pi(\omega)$ is not self-avoiding. Such walks must visit two distinct points $y,y^{\prime}$ in $\mathbb{Z}^{d}$} with $\pi(y)=\pi(y^{\prime})$. Diagrammatically, this is  [Uncaptioned image]  where the hollow circles are $0,x^{\prime}$ and the filled squares are $y,y^{\prime}$. By neglecting the mutual avoidance of the three diagram lines, the difference $\Gamma_{R,\beta}(x)-G_{R,\beta}(x)$ is therefore bounded above by
@@ V Further applications of the plateau theorem
The plateau theorem has been isolated here, and also applied to long-range models, for the first time. However, it has previously been used implicitly in multiple contexts. We now summarise the relevant literature; the conclusions are presented in Table 1. (Logarithmic corrections at the critical dimension are discussed in [30, 31, 35, 36, 37] for SR models.)
@@
The results concern the following symmetric kernels $J$ on $\mathbb{Z}^{d}$ and $\bar{J}_{R}$ on $\mathbb{T}_{R}^{d}$ for short-range and long-range models:
@@
  • spread-out LR: $J_{x,y}\asymp L^{-d}(1+|x-y|/L)^{-(d+\alpha)}$ with $\alpha\in(0,2)$ for $d>1$ and $\alpha\in(0,1)$ for $d=1$ ($L$ is fixed and large).
@@
  • For the SR $1$-component $\varphi^{4}$ model, (6)–(7) are proved in [40] for $d>4$. For the torus, a rigorous renormalisation group method has recently been used to prove that for $d>4$ the bounds of Theorem 1 hold exactly at the infinite-volume critical point for both $\chi_{R}(\beta_{c})$ and $G_{R,\beta_{c}}$ [37], and with logarithmic corrections for $d=4$. Stronger results have been proved on the hierarchical lattice for $d\geq 4$ [30, 31]. We discuss this in more detail in Section VI.
@@ VI Universal profiles
which extends the definition of $f_{n}$ to all n>2n>-2. The $n=0$ case reduces to $f_{0}(s)=I_{1}(s)$ (since $I_{3}(x)+sI_{1}(s)=1$) as in Table 2; this is the profile for SAW on the complete graph [30, Appendix B] and is consistent with SAW being the $n=0$ spin model [44]. The limiting case $f_{-2}(s)=s^{-1}$ is the correct profile for a Gaussian model. Percolation on the complete graph is the Erdős–Rényi random graph. Its profile function has been proved to be $f_{\mathrm{perc}}$ in [45, Corollary 3.9] (see [17, Appendix C] for the identification of the profile function with $f_{\mathrm{perc}}$). At the moment of final revision of our paper, equation (41) in Conjecture 4 has been proved for percolation on the high-dimensional torus [46, Theorem 2.1] for SR spread-out models in dimensions $d>6$, and for the nearest-neighbour model for $d$ much greater than $6$.
@@ VII Free boundary conditions
$\chi_{R}^{\mathrm{F}}({\beta_{\mathrm{c}}})\asymp R^{\alpha},\qquad G_{R,{\beta_{\mathrm{c}}}}^{\mathrm{F}}(0,x)\asymp\frac{1}{|x|^{d-\alpha}},$ (47)
with $x$ away from the box boundary for the latter. (Since translation invariance is broken, we write both points in $G_{R,{\beta_{\mathrm{c}}}}^{\mathrm{F}}(0,x)$; the origin is the centre of the box.) In contrast with (4) for PBC, there is no plateau in (47). This has been proved for the SR Ising model [47] (with the hypothesis supplied by [40, 22]) and SR percolation [48].
@@
Since $d>d_{\mathrm{c},\alpha}=\frac{\alpha}{2}d_{\mathrm{c}}$, the shift $R^{-\frac{1}{\nu}}=V^{-\frac{\alpha}{\gamma d}}$ is larger than the window width $V^{-\frac{2}{\gamma d_{\mathrm{c}}}}$, and the PBC and FBC windows do not overlap. This shifting of the PBC scaling behaviour to the FBC pseudocritical point was observed numerically for the 5D SR Ising model and is discussed in [9, 4] (contrary to the conclusions of [13]).
@@ Acknowledgements
We thank our collaborators on aspects of this work: Tom Hutchcroft, Emmanuel Michta, Romain Panis. The work of YL and GS was supported in part by NSERC of Canada. YL and GS thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme Stochastic systems for anomalous diffusion, where work on this paper was undertaken; this work was supported by EPSRC grant EP/R014604/1. The work of GS was supported in part by the Clay Mathematics Institute. JP thanks the Seoul National University for hospitality during a visit when this work was partially undertaken. JP was supported by Basic Science Research Program through the National Research Foundation of Korea funded by the Ministry of Science and ICT (RS2025-00518980).
@@ Appendix A Elementary lemmas
If the maximum in the lower summation limit is $1$ then $\xi/R\leq 1$ and ($5$) is of order $(\xi/R)^{a+\varepsilon}\leq(\xi/R)^{a}$. If instead the maximum is $\lceil 2\xi/R\rceil$ then (55) is bounded by a multiple of $(\xi/R)^{a+\varepsilon}(\xi/R)^{-\varepsilon}=(\xi/R)^{a}$. This completes the proof. ∎
@@ B.4 Ising model
For the SR $n$-component $|\varphi|^{4}$ model in the critical dimension $d=d_{\mathrm{c}}=4$, it is well-known and rigorously proved [53, 54] that $\gamma=2\nu=1$ with logarithmic exponents $\hat{\gamma}=2\hat{\nu}=\frac{n+2}{n+8}$. Also, the specific heat has exponent $\alpha=0$ for $n\geq 1$, with $\hat{\alpha}=\frac{4-n}{n+8}$ for $n=1,2,3$, whereas for $n=4$ the specific heat diverges as $\log\log t^{-1}$ and for $n>4$ it does not diverge at all [53]. In particular, $\hat{\alpha}$ is never strictly negative. Let $\hat{\theta}=\frac{4-n}{2(n+8)}=\frac{1}{2}-\hat{\gamma}$, for all $n\geq 1$ including $n\geq 4$. It is argued in [35, (3.6), (4.3)] that under PBC the finite-volume susceptibility has a log correction $\chi_{R}\asymp V^{1/2}(\log V)^{1/2}$ and that the window width is $V^{-1/2}(\log V)^{-\hat{\theta}}$. It was proved rigorously that $\chi_{R}(\beta_{c})\asymp V^{1/2}(\log V)^{1/2}$ in [37] using a rigorous RG analysis [55]. For the hierarchical model, precise asymptotics for both the torus susceptibility and window width were proved in [30].
At the high-temperature edge of the window, the infinite-volume correlation length therefore obeys
@@ Appendix D Renormalisation group
Universal ratios can be computed similarly. E.g., with the same $\nu=\nu_{\mathrm{c}}+sa_{d}^{-1}\bar{g}^{1/2}V^{-1/2}$, the moments of the total spin $\Phi_{R}=\sum_{x\in\mathbb{T}_{R}^{d}}\varphi_{x}$ obey, as $R\to\infty$,